From mar@roe.ac.uk Wed Mar 6 09:20:03 2013 Date: Sun, 3 Mar 2013 16:53:44 +0000 From: Mike Read To: VDFS Science Archive Team -- Clive Davenhall , Eckhard Sutorius , Mark Holliman , Nicholas Cross , Nigel Hambly , Rob Blake , Ross Collins , Stelios Voutsinas Cc: CCs for VDFS -- Alastair Edge , Andrew Lawrence , Bob Mann , Dante Minniti , James Dunlop , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Maria-Rosa Cioni , Matt Jarvis , Mike Irwin , Nigel Metcalfe , Norman Gray , Phil Lucas , Richard McMahon at IOA , Simon Dye , Stephen Warren , Tom Kerr , Tom Shanks Subject: WFAU VDFS Science Archives project team meeting minutes, 1st March 2013 Present: MAR, RPB, ETWS, RGM, NJC Apologies: JPE, AL, STV, MSH, NG, KTN, RGM, NCH, RSC DoNM: 10am Friday March 15th 2013 in the Villas meeting room NCH in the chair (room is booked). Actions discharged this time: ----------------------------- ACTION: RPB to provide cupcakes for the "Quarterly" meeting. Discharged, and very good they were too! Actions partly discharged but continuing: ----------------------------------------- ACTION: ETWS (with help from NJC as necessary) to add FeII 1.644 micron filter details into the WSA. Continues; all done apart from Vega-to-AB and zeropoint attributes which can now be done by NJC as sufficient data have been ingested Actions carried forward from 25/01/2013 meeting ----------------------------------------------- ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues; this one still has low priority. Specific points and new actions: -------------------------------- Project management: As mentioned at the "Quarterly" it was "Annual" PDR time again. RGM and AL would be conducting the interviews, forms to be issued shortly. WFCAM, VISTA and VST updates: Nothing to report Comments and issues arising from CASU minutes: No comments. Networking: Network working well and transfers are up-to-date. WSA/VSA/OSA/GES Operations: ETWS reported: "- Processing of VVV P89 catalogue data is nearly finished and ingests are half way through. - The latest VSA data for Nov and Dec 2012 have been transferred, JPEGs are created and pixel data is ingested." The recent CU4 updates have resulted in a major speed up for VVV ingests. However the current SQL view of vvvDetection, being made up of all the monthly tables, meant that there were sometimes collisions between ingest and release operations. It was suggested that an additional view could be created by setup programme that only referenced those monthly tables required for the current release. It was noted that we might not be using the latest CASU tools software. ACTION: NJC to contact CASU to enquire about the latest CASU software and the availability of a stacker for VST. We are expecting that the first GES data will be made available to us very shortly. Effort has been lined up to get the data ingested and released as soon as possible. Hardware and Systems: All working well. Survey Data Release: The VVV P87 release continues, currently still working through neighbour table and variability curation (CU6). VMC - currently in final release stage (CU19) and likely to finish over the weekend. VIKING and VHS P89 releases will follow the VMC. ACTION: MAR to do VHS QC (mainly deprecating duplicates) and contact Alastair about VIKING QC requirements. VIDEO - nothing from Matt but not really expecting new mosaics until around April. ATLAS - the P88 release database had been made however it was discovered that the colour attributes in atlasSource had not been populated. The latter stages of the release were being re-run and the release would be ready agian in a few days. RGM noted that the ATLAS release would be accessed via new user-interface which was nearly ready. ETWS mentioned that if would be useful if the new UI could be available under a test environment in addition to a live server so that changes made to say the schema browser could be checked before going live. MAR also asked if queries were logged by the new UI so that they could be readily accessed as this was often helpful. RGM requested that the OSA/ATLAS webpages be updated to accompany the release. MAR said that the content held in the old system had been updated recently and ETWS confirmed that the schema browser output was also current. ACTION: MAR to liaise with STV over updating new UI webpages. Software: ETWS reported: "- Worked with STV on the browser pages for the new UI. - Enhanced the code to allow smoother operations: a) easier check of all file shares to avoid running out of space; b) easier checks of the compilation status of our code base to avoid running outdated C++ code; c) enabled ingests by cuEventID and/or server name in addition to cuID. - After the test phase of running CU4 VVV ingests without immediate objID updates went well, I've move the update now completely into the FinaliseCatalogueIngest script." MAR reported that the image and crossID services had been modified so that they could be called by the new UI. Non-survey Data Release: ACTION: MAR to supply a list of non-survey projects to be released. Astrogrid deployment, VO & Data Analysis services: Nothing of note. Miscellaneous: Nothing of note. -- Scanned by iCritical.